Q1: Given a message “explanations”

a) Define a key, which is a combination of last names of all team members. For example, for a team of two persons ‘John Smith’ and ‘John Li’ the key will be ‘SmithLi’. You will use this key for encryption and decryption below

b) First encrypt the message using Vigenère cipher. Encrypt the resulting cipher-text using Playfair cipher.

c)Decrypt the ciphertext from part b above. Hint: First decrypt using Playfair cipher and then decrypt it again using Vigenère cipher
Answer #1

a) The random string of bits created to encrypt and decrypt the data or a message is the Key , In the given example the key is combination of last names Smith and Li that is SmithLi .

b) To encrypt the message "explanations" using key "SmithLi" by Vigenere cipher follow the steps:

Message : explanations

Key : SmithLi

Use the table given above , pair the first letter of message "e" with first letter of key "s" we get "w" as encryted letter , follow the same for second letter "x" pair with second letter of key "m" then the excrypted letter is j , by encrypting complete message we get " wjxehyiluwgz " encrypted message.

To further encrypt the message " wjxehyiluwgz" using playfair cipher with key SmithLi ,

The table used is :

S M I/J T H
L A B C D
E F G K N
O P Q R U
V W X Y Z

The message can be written as,

wj xe hy il uw gz

By using the rule of rectangle find the corresponding letters of w and j which form the rectangle in table , they are 'x' for 'w' an 'm' for 'j' , Hence find corresponding letters for all the pairs of message ,

The encrypted message is : xmvgtzsbpznx

c) To decrypt the message , first decrypt the message " xmvgtzsbpznx" using Playfair cipher ,

The table used is :

S M I T H
L A B C D
E F G K N
O P Q R U
V W X Y Z

Pair the letters of message as, xm vg tz sb pz nx

And follow the same process of encryption , the corresponding letters of xm is table are 'wi' like wise decrypt all the pairs to get " wixehyiluwgz " decrypted message .

The message obtained " wixehyiluwgz" from decryption is further decrypted using Vigenere cipher , To decrypt the message use the formula, p = c - k mod 26, where p is the plain text , c is the encoded text and k is the key = smithli,

For example : p = w-s mod 26 = 22- 18 = 4 ->e

By decoding using this formula the decoded message got is "ewplanations"